OpenStudy (anonymous):

Find the domain and range of a real function f(x) =sqrt 9- x^2 ?

OpenStudy (anonymous):

@satellite73

OpenStudy (zepp):

\[\sqrt{9-x^2}\]

OpenStudy (anonymous):

here we go again but this time lets work smarter

OpenStudy (zepp):

Do the same thing as @satellite73 showed you :D

OpenStudy (anonymous):

start with the equation of a circle with radius 3 and center at the origin. the equation for such a circle is \(x^2+y^2=9\) and i hope you have seen this before so lets solve for \(y\) and see what happens

OpenStudy (anonymous):

wat confused..... do this by that way

OpenStudy (anonymous):

\[x^2+y^2=9\] \[y^2=9-x^2\] \[y=\pm\sqrt{9-x^2}\] now this is not a function, because of the \(\pm\) but that is ok, get rid of the \(-\) part and get \(f(x)=\sqrt{9-x^2}\)

OpenStudy (anonymous):

so what you see is that this function is just the upper half of the circle, i.e. a semicircle, which radius 3 so the domain will be \([-3,3]\) and the range will bet \([0,3]\)

OpenStudy (anonymous):

@satellite73 any other way!!

OpenStudy (anonymous):



OpenStudy (anonymous):

ur answer is correct!

OpenStudy (anonymous):

yes, we can do just what we did before

OpenStudy (anonymous):

can u show that plzz

OpenStudy (experimentx):